What is the specific heat of lead if it takes 106 J to raise the temperature of a 75 g block by 11oC?
1
Expert's answer
2020-09-29T06:51:46-0400

Specific heat capacity is the amount of heat that will raise the temperature of 1kg of a substance by 1 kelvin. The SI unit of SHC is Joules per kilogram per kelvin(Jkg-1/K-1).

Heat=mass(kg)×specific heat capacity(c)×change in temperature

c=j/(mxt)

c=106÷(0.075×284.15)

c=106÷24.31

c=4.36Jkg-1 k-1